# Albrechtstr. 30, 49076 Osnabrück, Germany

### v1

@@ 3.4 Building Projectors from Autoencoder Embeddings
We continue to modify the standard projector by replacing its ReLU activation with SiLU (Sigmoid Linear Unit), sigmoid and . As with SimCLR, our modified projector is activated using the output features generated by its linear input layer. For this reason, the number of input features in the projector’s output layer must match the number of output features produced by the input layer. In order to avoid mismatches in layer dimensions during our experiments, we choose to programmatically apply these modifications to the projector’s architecture.
@@
3.5 End-to-End Training of Backbone and Projectors
In addition to the projection heads constructed in the previous section, we choose ResNet34 as our backbone over the much deeper ResNet50 architecture used in the original SimCLR paper. This choice is based on the fact, that the autoencoder implementation of Chadebec et al. [1] uses an embedding layer with 512 input features, matching the number of output features of ResNet34. ResNet variants deeper than ResNet34 increase their final feature dimensions from 512 to 2048 [5] and would therefore render all custom projection heads incompatible.
We train our modified SimCLR architecture as depicted in Figure 1 using mostly the same set of hyperparameters as proposed by Chen et al.[2]. In order to compensate for the shallower backbone, we choose to train for 150 epochs while using Stochastic Gradient Descent (SGD) as the optimizer with an initial learning rate of 30 % the batch size divided by 256 and momentum set to 0.9. We choose a weight decay of and dynamically adapt the learning rate using the Cosine Annealing scheduler with parameters and set to a fiftieth of the initial learning rate. We parameterize the NT-Xent loss using a fixed temperature value of . Training is conducted using minibatch sizes that allow for at least one full validation batch. For larger datasets like CIFAR10 and CIFAR100, we use a batch size of 1280 (see Table 1). All training runs use the same 70/10/20 dataset splits which were created in Section 3.3 to train the autoencoder embeddings. Individual dataset samples are augmented into a contrastive image pair using two randomly sampled transformations from SimCLR’s default transformation space .
@@ 3.5 End-to-End Training of Backbone and Projectors
Unlike during the training of the autoencoders, we now allow training runs to normalize datasets prior to augmentation using the standard score or -score method. We train and evaluate 144 different SimCLR models for each of the five image classification datasets with 72 different projection heads of which 36 include a pretrained autoencoder embedding. As part of these training runs, we also examine the performance impact of the following influencing factors:
Three latent dimensions (128, 256, 512) for the projector’s input layer.
Four different activations (ReLU, SiLU, sigmoid, ) of the projector.
Three latent dimensions (32, 64, 128) for the final projections.
-score normalization of datasets prior to training and evaluation.
3.6 Evaluating Learned Representations
For each trained backbone , we evaluate the quality of its contrastively learned representations by training a logistic regression model on the intermediate representations using a labeled test split. This approach follows the standard evaluation protocol of SimCLR and related methods. The regression model is represented by a single linear layer whose number of input features equals the dimensionality of . The number of its output features is dynamically determined by the the number of classes present in the test split. We train the logistic regression model for 50 epochs with no weight decay, a learning rate of and a minibatch size of 32. We apply standard softmax to the model’s logits to obtain class probabilities.
@@ 4 Results
We begin by summarizing the best test accuracies (Acc@1) achieved by our custom projection heads (ae) and the default 2-layer MLP (mlp) used in the SimCLR framework on five well-known image classification datasets (see Table 2). A first expected observation is that classification accuracies decrease with increasing number of classes, regardless of the type of projection head used. However, this correlation is observed to be strongly nonlinear, meaning the accuracies achieved on related datasets like CIFAR10 and CIFAR100, where the number of classes increases tenfold, worsen only by about a factor of three. Our absolute best result is recorded on the STL10 dataset where our custom projector outperforms the standard 2-layer MLP by 2.9 %. On average, we achieve an improvement in classification accuracy of 1.74% across all datasets.
| Dataset | Proj. Activation | Normalized | Acc@1 | |||||||
| Projector | ae | mlp | ae | mlp | ae | mlp | ae | mlp | ae | mlp |
| Imagenette | 256 | 256 | Sigmoid | Sigmoid | 128 | 128 | False | False | 89.0 % | 88.7 % |
| STL10 | 256 | 256 | Sigmoid | Sigmoid | 128 | 128 | False | True | 72.3 % | 69.4 % |
| CIFAR10 | 128 | 256 | Tanh | Sigmoid | 128 | 64 | False | False | 66.5 % | 64.9 % |
| FGVCAircraft | 256 | 256 | ReLU | ReLU | 128 | 32 | True | False | 25.7 % | 24.0 % |
| CIFAR100 | 512 | 512 | SiLU | ReLU | 64 | 64 | False | False | 23.6 % | 21.5 % |
@@
Table 2 also summarizes the best evaluation results obtained after altering the values of the four major influencing factors listed in Section 3.5. We choose to denote the number of output features in the input layer of the projector and to represent the dimensionality of the final projections, i. e. the number of output features of the projector. Our results listed in Table 2 indicate, that in order to achieve peak classification accuracy across all datasets, the number of output features in the output layer of our custom projector has to be increased in some cases when compared to the default 2-layer MLP. On average, however, we see a significant reduction in projector width when using a pretrained input layer as outlined in Table 4.
@@
We also observe, that activating the projector using either the sigmoid or function outperforms ReLU on all datasets with ten classes, whereas more complex datasets like FGVCAircraft and CIFAR100 work better with linear units. Counterintuitively, we find -score normalization to reduce peak classification accuracy for both projectors across a majority of different datasets.
@@
In order to further analyze the performance of the different projection heads, we report key statistical indicators for the distribution of test accuracies across all runs grouped by projector type and dataset. The upper half of Table 3 shows the statistics for the training runs in which -score normalization was performed, whereas the lower half lists the results without dataset normalization. Statistical analysis reveals, that by using our custom projector the interquartile range (midspread) of test accuracies is almost always reduced, indicating a general improvement in model stability. This observation can also be made for datasets where the absolute range and standard deviation of test accuracies is larger, hinting at an increased number of outliers when using the ae projector. In addition to the maximum values, the mean result also suggest significant improvements in average classification accuracy when using our pretrained projectors over the default 2-layer MLP. We find this observation to be true for all datasets, even the ones with 30 and 100 classes.
@@
| Projector | ae | mlp | ae | mlp | ae | mlp | ae | mlp | ae | mlp |
| minimum | 41.9 % | 68.4 % | 42.6 % | 44.4 % | 56.8 % | 44.6 % | 20.4 % | 20.1 % | 8.3 % | 3.2 % |
| maximum | 88.5 % | 88.5 % | 70.8 % | 69.4 % | 66.3 % | 64.1 % | 25.7 % | 23.9 % | 23.3 % | 21.5 % |
| average | 83.6 % | 83.1 % | 64.4 % | 59.7 % | 62.6 % | 58.5 % | 23.4 % | 22.1 % | 19.0 % | 16.2 % |
| range | 46.6 % | 20.1 % | 28.2 % | 24.9 % | 9.4 % | 19.5 % | 5.3 % | 3.8 % | 15.0 % | 18.3 % |
| midspread | 3.2 % | 3.3 % | 3.8 % | 6.6 % | 3.3 % | 3.8 % | 1.7 % | 1.0 % | 3.1 % | 4.7 % |
| std. dev. | 8.0 % | 4.3 % | 6.8 % | 5.6 % | 2.4 % | 3.6 % | 1.2 % | 0.9 % | 3.1 % | 4.2 % |
| minimum | 66.9 % | 62.5 % | 39.1 % | 32.3 % | 44.4 % | 48.8 % | 20.9 % | 20.6 % | 5.6 % | 3.2 % |
| maximum | 89.0 % | 88.7 % | 72.3 % | 67.8 % | 66.5 % | 64.9 % | 25.2 % | 24.0 % | 23.6 % | 21.5 % |
| average | 85.0 % | 82.8 % | 65.1 % | 58.8 % | 61.3 % | 58.7 % | 23.0 % | 22.3 % | 19.4 % | 16.6 % |
@@
| midspread | 2.2 % | 3.2 % | 4.8 % | 10.2 % | 2.9 % | 5.5 % | 1.4 % | 1.1 % | 2.6 % | 4.0 % |
| std. dev. | 3.8 % | 6.1 % | 6.0 % | 8.0 % | 4.4 % | 4.0 % | 1.0 % | 0.8 % | 3.1 % | 4.1 % |
In Table 4 we examine the influence of varying the width of the projector by changing the number of output features in the input and output layers. We deliberately constrain our experiments to network width rather than depth to avoid any major architectural changes compared to the default 2-layer MLP. The average test accuracies listed in the upper half of Table 4 are obtained by varying the number of output features in the projector’s output layer, whereas the bottom half contains the average results achieved by changing the number of output features in the projector’s output layer. We monotonically decrease the output features in the input layer from 512 to 128 and in the output layer from 128 to 32 in order to generate contraction within the projector. This approach is consistent with the standard SimCLR approach, but instead of limiting the final embeddings to 128 feature dimensions, it allows for even smaller projection spaces. Our experiments indicate, that using a pretrained autoencoder embedding inside the projector can reliably decrease the number of output features needed in both layers while achieving better classification results. On average, ae training runs using 128 or 256 as and 32 or 64 for achieve higher test accuracies on all datasets compared to their wider mlp counterparts.
@@
In Table 5 we summarize the average test accuracies as a function of projector activation. The results show, that on average, our modified ae projector prefers the classical activation function over modern linear units like ReLU and SiLU. We find this observation to be true for almost all datasets, although for reaching peak performance on Imagenette and STL10, sigmoid is still preferred (see Table 2). It should be noted that on average, the mlp projector also works best with the sigmoid activation function when trained on the Imagenette, CIFAR10 and FGVCAircraft datasets.
| Projector | ae | mlp | ae | mlp | ae | mlp | ae | mlp | ae | mlp |
| ReLU | 85.0 % | 84.0 % | 64.5 % | 59.9 % | 62.9 % | 59.3 % | 23.0 % | 22.0 % | 20.1 % | 19.3 % |
| SiLU | 85.3 % | 84.5 % | 64.7 % | 59.7 % | 61.1 % | 59.1 % | 23.2 % | 22.1 % | 19.6 % | 18.6 % |
| Sigmoid | 81.8 % | 84.7 % | 64.6 % | 58.5 % | 61.2 % | 59.4 % | 23.1 % | 22.8 % | 16.6 % | 13.4 % |
| Tanh | 85.3 % | 78.4 % | 65.3 % | 58.9 % | 62.6 % | 56.6 % | 23.4 % | 22.0 % | 20.4 % | 14.3 % |
@@
In Table 6 we finally report our initial test results comparing the performance impact of frozen and non-frozen autoencoder embeddings. In the non-frozen setting, the weights within each pretrained embedding layer are allowed to change in the course of SimCLR’s end-to-end training. Separate training runs, in which we use sigmoid as the projector’s activation function and do not apply -score normalization to the datasets, clearly show an increase in peak and average test accuracy after freezing the weights within the pretrained autoencoder embedding. We argue that these results justify the decision to carry out all ae training runs using only frozen embeddings.
| Acc@1 | max. | avg. | max. | avg. | max. | avg. | max. | avg. | max. | avg. |
| Frozen | 89.0 % | 85.2 % | 72.3 % | 64.1 % | 65.4 % | 59.8 % | 25.2 % | 23.3 % | 22.6 % | 17.5 % |
| Not frozen | 87.9 % | 86.2 % | 71.3 % | 60.5 % | 63.7 % | 56.4 % | 24.3 % | 23.1 % | 21.9 % | 17.1 % |
5 Discussion
@@
Our experiments demonstrate that the SimCLR framework can benefit from using pretrained autoencoder embeddings over a randomly initialized input layer to its nonlinear projection head. This finding does therefore confirm our initial hypothesis outlined in Section 1. We explain the success of our approach by the fact that pretraining even a simple autoencoder on a given dataset can capture meaningful representations in its embedding layer. These features can then be exploited within the SimCLR framework by offering a priori knowledge about the high-level structure of the training data. By including these predetermined features into the projector, subsequent training iterations can benefit immediately from these representations. This effect cannot be achieved using the standard 2-layer MLP as the weights of its input layer are randomly initialized. Due to the NT-Xent loss being applied only to the projections , it is reasonable to assume that using meaningful high-level representations within the projector can guide gradient descent more effectively, especially in the early stages of training.
@@
Considering the architectural simplicity of our autoencoder, we find the absolute and average improvements in classification accuracy to be satisfactory, especially in the case of datasets containing a large number of classes. It should also be noted, that in our experiments we neither adapt the architecture of the backbone nor the projection head to compensate for such datasets or different sample resolutions. Regardless of dataset complexity, we assume absolute improvements in classification accuracy to be achievable through deeper and wider backbones trained for longer periods of time. We also suspect that pretraining more advanced autoencoder architectures could further increase the projector’s overall performance.
@@
Considering the main purpose of autoencoders, we conjecture that the same features which are suitable for image reconstruction tasks, can also enhance contrastive representation learning as demonstrated in this paper. Additional benefits of our approach can be observed in architectural efficiency based on the decreased number of output features required for the pretrained projector to surpass the standard MLP in average classification accuracy (see Table 4). This finding suggests that the implicit training of the standard MLP projector within the SimCLR framework is often less efficient than the independent pretraining of an autoencoder embedding.
@@ 6 Conclusions
Increase in accuracy: Using a pretrained and frozen autoencoder embedding in SimCLR’s default projector can increase classification accuracy by up to 2.9 % or 1.7 % on average (see Table 2).
Architectural efficiency: On average, a pretrained autoencoder embedding reliably decreases projector width and the dimensionality of the associated projection space (see Table 4).
Projector stability: By freezing the weights within the pretrained embeddings, we demonstrate the long lasting efficacy of our projector across multiple training epochs. If we allow the weights to be changed during training, a noticeable decrease in test accuracy is observed (see Table 6).
@@
[…]
The authors have no competing interests to declare that are relevant to the content of this article.

### v2

@@ IV-B Building projectors from autoencoder embeddings
We continue to modify the standard projector by replacing its ReLU activation with SiLU (Sigmoid Linear Unit) [Elfwing2018], sigmoid and . As with SimCLR, our modified projector is activated using the output features generated by its linear input layer. For this reason, the number of input features in the projector’s output layer must match the number of output features produced by the input layer. In order to avoid mismatches in layer dimensions during our experiments, we choose to programmatically apply these modifications to the projector’s architecture.
@@
IV-C End-to-end training of backbone and projectors
In addition to the projection heads constructed in the previous section, we choose ResNet34 as our backbone over the much deeper ResNet50 architecture used in the original SimCLR paper. This choice is based on the fact, that the autoencoder implementation of Chadebec et al. [Chadebec2022] uses an embedding layer with 512 input features, matching the number of output features of ResNet34. ResNet variants deeper than ResNet34 increase their final feature dimensions from 512 to 2048 [He2016] and would therefore render all custom projection heads incompatible.
We train our modified SimCLR architecture as depicted in Fig. 1 using mostly the same set of hyperparameters as proposed by Chen et al. [Chen2020]. In order to compensate for the shallower backbone, we choose to train for 150 epochs while using Stochastic Gradient Descent (SGD) as the optimizer with an initial learning rate of 30 % the batch size divided by 256 and momentum set to 0.9. We choose a weight decay of and dynamically adapt the learning rate using the Cosine Annealing scheduler [Loshchilov2017] with parameters and set to a fiftieth of the initial learning rate. We parameterize the NT-Xent loss using a fixed temperature value of . Training is conducted using minibatch sizes that allow for at least one full validation batch. For larger datasets like CIFAR10 and CIFAR100, we use a batch size of 1280 (see Table I). All training runs use the same 70/10/20 dataset splits, c. f., Section IV-A, to train the autoencoder embeddings. Individual dataset samples are augmented into a contrastive image pair using two randomly sampled transformations from SimCLR’s default transformation space .
@@ IV-C End-to-end training of backbone and projectors
We train and evaluate 144 different SimCLR models for each of the five image classification datasets with 72 different projection heads of which 36 include a pretrained autoencoder embedding. As part of these training runs, we also examine the performance impact of the following influencing factors:
Three projector latent dimensions (128, 256, 512)
Four projector activations (ReLU, SiLU, sigmoid, )
Three projection latent dimensions (32, 64, 128)
IV-D Evaluating learned representations
For each trained backbone , we evaluate the quality of its contrastively learned representations by training a logistic regression model on the intermediate representations using a labeled test split. This approach follows the standard evaluation protocol of SimCLR and respective related methods. The regression model is represented by a single linear layer whose number of input features equals the dimensionality of . The number of its output features is dynamically determined by the the number of classes present in the test split. We train the logistic regression model for 50 epochs with no weight decay, a learning rate of and a minibatch size of 32. We apply standard softmax in order to the model’s logits to obtain class probabilities.
@@ V Results
We begin by summarizing the best test accuracies (Acc@1) achieved by our custom projection heads (ae) and the default 2-layer MLP (mlp) used in the SimCLR framework on five well-known image classification datasets (see Table II). A first expected observation is that classification accuracies decrease with increasing number of classes, regardless of the type of projection head used. However, this correlation is observed to be strongly nonlinear, meaning the accuracies achieved on related datasets like CIFAR10 and CIFAR100, where the number of classes increases tenfold, worsen only by about a factor of three. Our absolute best result is recorded on the STL10 dataset where our custom projector outperforms the standard 2-layer MLP by 2.9 %. On average, we achieve an improvement in classification accuracy of 1.74 % across all datasets.
| Dataset | Proj. Activation | Acc@1 | ||||||
| Head | ae | mlp | ae | mlp | ae | mlp | ae | mlp |
| Imagenette | 256 | 256 | Sigmoid | Sigmoid | 128 | 128 | 89.0 % | 88.7 % |
| STL10 | 256 | 256 | Sigmoid | Sigmoid | 128 | 128 | 72.3 % | 69.4 % |
| CIFAR10 | 128 | 256 | Tanh | Sigmoid | 128 | 64 | 66.5 % | 64.9 % |
| FGVCAircraft | 256 | 256 | ReLU | ReLU | 128 | 32 | 25.7 % | 24.0 % |
| CIFAR100 | 512 | 512 | SiLU | ReLU | 64 | 64 | 23.6 % | 21.5 % |
@@
Table II also summarizes the best evaluation results obtained after altering the values of the four major influencing factors listed in Section IV-C. We choose to denote the number of output features in the input layer of the projector and to represent the dimensionality of the final projections, i. e. the number of output features of the projector. Our results listed in Table II indicate, that in order to achieve peak classification accuracy across all datasets, the number of output features in the output layer of our custom projector has to be increased in some cases when compared to the default 2-layer MLP. On average, however, we see a significant reduction in projector width when using a pretrained input layer as outlined in Table IV.
@@
We also observe, that activating the projector using either the sigmoid or function outperforms ReLU on all datasets with ten classes, whereas more complex datasets like FGVCAircraft and CIFAR100 work better with linear units.
@@
In order to further analyze the performance of the different projection heads, we report key statistical indicators for the distribution of test accuracies across all runs grouped by projector type and dataset (see Table III). Statistical analysis reveals, that by using our custom projector the interquartile range (midspread) of test accuracies is almost always reduced, indicating a general improvement in model stability. This observation can also be made for datasets where the absolute range and standard deviation of test accuracies is larger, hinting at an increased number of outliers when using the ae projector. In addition to the maximum values, the mean result also suggest significant improvements in average classification accuracy when using our pretrained projectors over the default 2-layer MLP. We find this observation to be true for all datasets, even the ones with 30 and 100 classes.
@@
| Head | ae | mlp | ae | mlp | ae | mlp | ae | mlp | ae | mlp |
| min. | 66.9 % | 62.5 % | 39.1 % | 32.3 % | 44.4 % | 48.8 % | 20.9 % | 20.6 % | 5.6 % | 3.2 % |
| max. | 89.0 % | 88.7 % | 72.3 % | 67.8 % | 66.5 % | 64.9 % | 25.2 % | 24.0 % | 23.6 % | 21.5 % |
| avg. | 85.0 % | 82.8 % | 65.1 % | 58.8 % | 61.3 % | 58.7 % | 23.0 % | 22.3 % | 19.4 % | 16.6 % |
@@
| IQR | 2.2 % | 3.2 % | 4.8 % | 10.2 % | 2.9 % | 5.5 % | 1.4 % | 1.1 % | 2.6 % | 4.0 % |
| SD | 3.8 % | 6.1 % | 6.0 % | 8.0 % | 4.4 % | 4.0 % | 1.0 % | 0.8 % | 3.1 % | 4.1 % |
In Table IV, we examine the influence of varying the width of the projector by changing the number of output features in the input and output layers. We deliberately constrain our experiments to network width rather than depth to avoid any major architectural changes compared to the default 2-layer MLP. The average test accuracies listed in the upper half of Table IV are obtained by varying the number of output features in the projector’s input layer, whereas the bottom half contains the average results achieved by changing the number of output features in the projector’s output layer. We monotonically decrease the output features in the input layer from 512 to 128 and in the output layer from 128 to 32 in order to generate contraction within the projector. This approach is consistent with the standard SimCLR approach, but instead of limiting the final embeddings to 128 feature dimensions, it allows for even smaller projection spaces. Our experiments indicate, that using a pretrained autoencoder embedding inside the projector can reliably decrease the number of output features needed in both layers while achieving better classification results. On average, ae training runs using 128 or 256 as and 32 or 64 for achieve higher test accuracies on all datasets compared to their wider mlp counterparts.
(top: , bottom: )
@@
In Table V, we summarize the average test accuracies as a function of projector activation. The results show, that on average, our modified ae projector prefers the classical activation function over modern linear units like ReLU and SiLU. We find this observation to be true for almost all datasets, although for reaching peak performance on Imagenette and STL10, sigmoid is still preferred (see Table II). It should be noted that on average, the mlp projector also works best with the sigmoid activation function when trained on the Imagenette, CIFAR10 and FGVCAircraft datasets.
| Avg. Acc@1 | ReLU | SiLU | Sigmoid | Tanh | ||||
| Head | ae | mlp | ae | mlp | ae | mlp | ae | mlp |
| Imagenette | 85.0 % | 84.0 % | 85.3 % | 84.5 % | 81.8 % | 84.7 % | 85.3 % | 78.4 % |
| STL10 | 64.5 % | 59.9 % | 64.7 % | 59.7 % | 64.6 % | 58.5 % | 65.3 % | 58.9 % |
| CIFAR10 | 62.9 % | 59.3 % | 61.1 % | 59.1 % | 61.2 % | 59.4 % | 62.6 % | 56.6 % |
| FGVCAircraft | 23.0 % | 22.0 % | 23.2 % | 22.1 % | 23.1 % | 22.8 % | 23.4 % | 22.0 % |
| CIFAR100 | 20.1 % | 19.3 % | 19.6 % | 18.6 % | 16.6 % | 13.4 % | 20.4 % | 14.3 % |
@@
In Table VI, we finally report our initial test results comparing the performance impact of frozen and non-frozen autoencoder embeddings. In the non-frozen setting, the weights within each pretrained embedding layer are allowed to change in the course of SimCLR’s end-to-end training. Separate training runs, in which we use sigmoid as the projector’s activation function, clearly show an increase in peak and average test accuracy after freezing the weights within the pretrained autoencoder embedding. We argue that these results justify the decision to carry out all ae training runs using only frozen embeddings.
| Acc@1 | Frozen | Not frozen | ||
| max. | avg. | max. | avg. | |
| Imagenette | 89.0 % | 85.2 % | 87.9% | 86.2% |
| STL10 | 72.3 % | 64.1 % | 71.3 % | 60.5 % |
| CIFAR10 | 65.4 % | 59.8 % | 63.7 % | 56.4 % |
| FGVCAircraft | 25.2 % | 23.3 % | 24.3 % | 23.1 % |
| CIFAR100 | 22.6 % | 17.5 % | 21.9 % | 17.1 % |
VI Discussion
@@
Our experiments demonstrate that the SimCLR framework can benefit from using pretrained autoencoder embeddings over a randomly initialized input layer to its nonlinear projection head. This finding does therefore confirm our initial hypothesis outlined in Section I. We explain the successful application of our approach by the fact that pretraining even a simple autoencoder on a given dataset can capture meaningful representations in its embedding layer. These features can then be exploited within the SimCLR framework by offering a priori knowledge about the high-level structure of the training data. By including these predetermined features into the projector, subsequent training iterations can benefit immediately from these representations. This effect cannot be achieved using the standard 2-layer MLP as the weights of its input layer are randomly initialized.
Due to the NT-Xent loss being applied only to the projections , it is reasonable to assume that using meaningful high-level representations within the projector can guide gradient descent more effectively, especially in the early stages of training.
@@
Given the architectural simplicity of our autoencoder, we find the absolute and average improvements in classification accuracy to be satisfactory, especially in the case of datasets containing a large number of classes. It should also be noted, that in our experiments we neither adapt the architecture of the backbone nor the projection head to compensate for such datasets or different sample resolutions.
Regardless of dataset complexity, we assume absolute improvements in classification accuracy to be achievable through deeper and wider backbones trained for longer periods of time. We also suspect that pretraining more advanced autoencoder architectures could further increase the projector’s overall performance.
@@
Considering the main purpose of autoencoders, we conjecture that the same features which are suitable for image reconstruction tasks, can also enhance contrastive representation learning as demonstrated in this paper. Additional benefits of our approach can be observed in architectural efficiency based on the decreased number of output features required for the pretrained projector to surpass the standard MLP in average classification accuracy (see Table IV). This finding suggests that the implicit training of the standard MLP projector within the SimCLR framework is often less efficient than the independent pretraining of an autoencoder embedding.
@@ VII Conclusions
Increase in accuracy: Using a pretrained and frozen autoencoder embedding in SimCLR’s default projector can increase classification accuracy by up to 2.9 % or 1.7 % on average (see Table II).
Architectural efficiency: On average, a pretrained autoencoder embedding reliably decreases projector width and the dimensionality of the associated projection space (see Table IV).
Projector stability: By freezing the weights within the pretrained embeddings, we demonstrate the long lasting efficacy of our projector across multiple training epochs. If we allow the weights to be changed during training, a noticeable decrease in test accuracy is observed (see Table VI).